OpenStudy (anonymous):

Which of the following is equivalent to tanxcotx? A)cos^2x/ sin^2x b) 1 C) cos^2x/sin^2x d) cscx i got b

OpenStudy (ash2326):

@jaquise2013 \[\tan x=\frac{\sin x}{\cos x}\] \[\cot x=\frac{\cos x}{\sin x}\] \[\tan x\times \cot x =\frac{\sin x}{\cos x}\times \frac{\cos x}{\sin x}=???\]